# Error Correction in Dynamical Codes

### v1

@@ 3.1 Definition of Distance for a Dynamical Code
[…]
@@
We begin by outlining the problem under consideration. For a given dynamical code, what is the set of errors that an ISG can correct given the syndrome information acquired through subsequent rounds of measurements? Using this information, what is the distance of the ISG? Without loss of generality, we examine the ISG at round 0 of the code and assume that all syndromes for this ISG come from the measurement sequence. To simplify the problem, we assume that the errors only occur at round 0 of the code and we also assume that if the errors are trivial, we know what the syndromes for the stabilizers in round 0 will be.
@@
Another example is the subsystem surface codes with three-qubit check operators in [28]. In this case, $d_{\mathrm{subsystem}}\leq d_{\mathrm{ISG}}$. If the gauges are fixed in the x basis, then the distance for a $[[3L^{2},2,L]]$ code is given by $L$ for $X$ logical and $2L$ for $Z$ logical, but for $d_{\mathrm{subsystem}}$, the minimum weight of dressed logical is $L$ for both $X$ and $Z$ logical operators.
@@ 3.2 Masking and Unmasked Distance of a Dynamical Code
One can extend the definition of $T$ to that of a group: Define $\tilde{T}=\langle T\rangle U$. Then, $\tilde{T}$ consists of the group of unmasked and temporarily masked stabilizers. We have $U\subseteq\tilde{T}$ and $P=\langle S\rangle\backslash\tilde{T}$.
@@
However, there are cases where even though two errors differ on the syndrome for masked generators, they do not affect our measurements of logical states. Let’s assume $E^{\dagger}F\in\mathcal{N}(U)\backslash\mathcal{N}(S)$, so $E$ and $F$ have the same unmasked syndrome but different masked syndromes. If $L$ and $L^{\prime}$ are two representatives for the same logical qubit, then measuring either is normally a valid way of measuring the logical value. Now, if $[E^{\dagger}F,L]=0$, then it does not matter whether we have an error $E$ or an error $F$; in either case, we will get the same eigenvalue for $L$ when we measure it. On the other hand, if $\{E^{\dagger}F,L^{\prime}\}=0$, then the error $E$ followed by measuring $L^{\prime}$ will give us a different result than the error $F$ followed by measuring $L^{\prime}$. In other words, we can pick a coset of logical representatives for which $[E^{\dagger}F,L]=0$ and this will allow us to correctly deduce the logical outcomes despite a larger set of errors that now consists of indistinguishable errors like $E$ and $F$.
@@
The distance is defined as the minimum weight of $E^{\dagger}F$ that does not satisfy Eq. (8), which is the complement of the set $[(P_{\mathrm{n}}-\mathcal{N}(U))\cup\langle S\rangle]\cdot\mathcal{G}$. This is exactly given by
@@
With the fixed choices of destabilizers, we have completely determined the set of bare logical operators, up to $U$. We see that there are 3 equivalent notions here: Choosing a gauge group $\mathcal{G}$ is the same as fixing the set of destabilizers which is also equivalent to choosing the bare logical representatives. We can see that doing so allows us to correct for more errors, since it also allows for errors that are destabilizers of the masked stabilizers, because these errors satisfy the condition that $[E^{\dagger}F,L]=0$, for all bare logical representatives $L$s of the new gauge group.
@@ 3.3 Logical Operators of a Dynamical Code
Suppose we initialize a dynamical code that consists of $l$ rounds of measurements in an eigenstate of $L_{0}$, which is fixed as our choice of logical representative for the $0^{\mathrm{th}}$ ISG, with the notations illustrated in Figure 3. Each round of measurements consists of commuting Pauli operators which form the set $M_{i}$ for the $i^{\mathrm{th}}$ round. Let $E_{k}$ be the net error from the time after the $k^{\mathrm{th}}$ round of measurements till the final round of the code. Then, $E_{0}$ is the net error since initialization till the final round of the code. The Pauli operators $L_{0}$ and $L_{l}$ are related by $L_{l}=L_{0}\cdot s_{0}\cdot\prod_{i=1}^{l-1}m_{i}$ up to ±$\pm 1$ sign difference. $m_{i}$ is an element of $\langle M_{i}\rangle$ and $O(m_{i})$ is the outcome of $m_{i}$ which can be obtained through the outcomes of elements in $M_{i}$. $s_{0}$ is an element of $\langle S_{0}\rangle$ which is the stabilizer group for round 00 of the dynamical code and $O(s_{0})$ is the outcome of $s_{0}$. $A\odot B$ is the binary symplectic product of A with B as defined in Subsection 2.3.
@@
The case where errors only occurred in round 00 follows as a special case.
@@
Suppose in round 1, we have a logical operator $L_{1}$, stabilizer $s_{1}$ in $S$ and error $e$ that occurred before round 1. Let $e\ket{\psi}$ be the stabilizer state in round 1 after $e$ occurs. In round 2, the logical operator remains as $L_{1}$, $s_{2}$ is a new stabilizer that anti-commutes with $e$ but commutes with $s_{1}$ and in round 3, we measure $s_{3}$ that anti-commutes with both $L_{1}$ and $s_{2}$ so that the new logical in rounds 3 is now updated to $L_{1}\cdot s_{2}$ and we have $s_{3}$ in the stabilizer group. This is illustrated in the table below:
ISG
Stabilizers
Logical
Events
$1$
$s_{1},\cdots$
$L_{1}$
@@
$3$
$s_{1},s_{3},\cdots$
$L_{1}\cdot s_{2}$
Assuming $[L_{1},e]=0$, even though $L_{1}\cdot s_{2}$ anti-commutes with $e$, the correct outcome of the logical operator is just given by $=O(L_{1})\cdot O(s_{2})$, independent of the error $e$ that happened earlier in the measurement schedule.
$(L_{1}\cdot s_{2})\left(1\pm s_{3}\right)\left(1\pm s_{2}\right)e\ket{\psi}$
$=$ $\left(1\pm s_{3}\right)(L_{1}\cdot s_{2})\left(1\pm s_{2}\right)e\ket{\psi}$
$=$ $\left(1\pm s_{3}\right)(\pm 1)\left(1\pm s_{2}\right)L_{1}\cdot e\ket{\psi}$
$=$ $O(s_{2})\cdot O(L_{1})\left(1\pm s_{3}\right)\left(1\pm s_{2}\right)e\ket{\psi}\>(\mathrm{assuming\>}[L_{1},e]=0)$
@@
On the other hand, if the error $e$ occurred in round 2, after measuring $s_{2}$, we see a change in the logical outcome. Here, we again assume that $[L_{1},e]=0$
ISG
Stabilizers
Logical
Events
$1$
$s_{1},\cdots$
$L_{1}$
@@
$3$
$s_{1},s_{3},\cdots$
$L_{1}\cdot s_{2}$
$(L_{1}\cdot s_{2})\left(1\pm s_{3}\right)e\left(1\pm s_{2}\right)\ket{\psi}$
$=$ $\left(1\pm s_{3}\right)e(-1)(L_{1}\cdot s_{2})\left(1\pm s_{2}\right)\ket{\psi}$
$=$ $\left(1\pm s_{3}\right)e(-1)(\pm 1)\left(1\pm s_{2}\right)L_{1}\ket{\psi}$
$=$ $-1\cdot O(s_{2})\cdot O(L_{1})\left(1\pm s_{3}\right)e\left(1\pm s_{2}\right)\ket{\psi}\>(\mathrm{assuming\>}[L_{1},e]=0)$
@@ 3.4 Syndromes through Measurements in a Dynamical Code
[…]
@@
As the decoding problem is much harder for the situation where errors occur in all rounds of the dynamical code and can depend on the details of the dynamical code, we consider only the simpler scenario with errors happening only in round 00 of the dynamical code, so that the outcomes for stabilizers in round 0 correspond only to syndromes of errors in round 0. From here, we can obtain an upper bound in the distance of the code.
@@ 4 The Distance Algorithm
We consider a dynamical code $\mathcal{Q}$ with $l$ rounds of measurements. The ISGs are denoted by $S_{0},S_{1},\cdots,S_{l}$ starting with the ISG $S_{0}$ in round 00. The sets of commuting measurements in each round are denoted by $M_{1},M_{2},\cdots$, where $M_{i}$ is the set of measurements made after the $(i-1)^{\mathrm{th}}$ ISG.
The $l$ rounds of measurements can also be interpreted as the window of measurements for obtaining the syndrome information to decode for errors in round 00. We assume that some errors occurred in round 00 and no more errors happen in the later rounds of measurements and we only use $l$ rounds of measurements as our measurement window for learning the syndrome bits in round 00. To gain a better understanding of the distance of the ISG in round 00, we need to know for each syndrome bit in round 00, which of the following 3 cases does it fall under given only the first $l$ rounds of measurements:
@@ 4.3 The Distance Algorithm
      Case 1: $[c,m]=[v,m]=0,\forall c\in C,\forall v\in V$ Update rule: Add $m$ to $V$ if $m\notin\langle V\rangle$ and record down $O(m)$.
@@
      Case 3: $[c,m]=\{v,m\}=0,\forall c\in C$ and for some $v\in V$
@@
      Update rule: Add $m$ to $V$ and record down $O(m)$ for $m$. Remove an element $m_{j}$ in $V$ where $m_{j}$ anti-commutes with $m$. Then, for all elements $m_{i}\in V$ that anti-commute with $m$, do the following update: $m_{i}\to m_{i}\cdot m_{j}$ with $O(m_{i})\cdot O(m_{j})$ as its outcome. For all elements $c_{i}\in C$ that anti-commute with $m$, do the following update: $c_{i}\to c_{i}\cdot m_{j}$. If $s_{i}\in\langle S_{0}\rangle$ is the stabilizer associated with $c_{i}$, then the stabilizer associated with $c_{i}\cdot m_{j}$ is $s_{i}$. If the operator and its outcome for $c_{i}\cdot s_{i}$ is given by $\tilde{m},O(\tilde{m})$ respectively, then for $c_{i}\cdot m_{j}$, the operator $\tilde{m}\cdot m_{j}$ has an outcome given by $O(\tilde{m})\cdot O(m_{j})$.
@@ 4.3.1 Unmasked Stabilizers and their Syndromes
Here, $[\alpha_{i},\beta_{i}]$ is a non-zero 2n-dimensional vector and $\bullet$ stands for an arbitrary n-dimensional vector.
The generators of $\langle C\rangle\cap\langle V\rangle$ are given by the set $\{F^{-1}([e_{i},f_{i}])\},1\leq i\leq j$ and the identities from the 00 vector rows of Eq. (17), as each 00 vector corresponds to a different associated stabilizer generator.
@@ 4.3.2 Permanently Masked Stabilizers and their Destabilizers
    Initialize $L$ as an empty set to keep track of the gauge operators for $\mathcal{Q}_{2}$
@@ 4.4 Examples for the Distance Algorithm
Round
$C$
$V$
00
$x_{1}x_{2}x_{3}x_{4}x_{5}x_{6}$
$1$
$x_{1}x_{2}x_{3}x_{4}x_{5}x_{6}$
$x_{5}x_{6}$
$2$
$x_{1}x_{2}x_{3}x_{4}$
$z_{6}z_{7}$
$3$
$x_{1}x_{2}x_{3}x_{4}$
$z_{6}z_{7},x_{1}x_{2},x_{3}x_{4}$
$\langle C\rangle\cap\langle V\rangle=\{x_{1}x_{2}x_{3}x_{4}\}$ Using algorithm 2, we can substitute into the equation $O(s_{j})=O(u_{j})\cdot O(u_{j}\cdot s_{j})$ with $O(u_{j})=O(x_{1}x_{2}x_{3}x_{4})$ and $O(u_{j}\cdot s_{j})=O(x_{5}x_{6})$. This gives the outcome for $x_{1}x_{2}x_{3}x_{4}x_{5}x_{6}$ which is an unmasked stabilizer for this measurement sequence.
@@ 4.5 Proof for the Distance Algorithm
Let $u$ be an element in $\tilde{U}$ and let its associated stabilizer be $s_{u}$. We prove that both $O(u)$ and $O(u\cdot s_{u})$ are known. This shows that $O(s_{u})$ has a known syndrome given by $O(u)\cdot O(u\cdot s_{u})$. Since $u\in\langle V\rangle$, $O(u)$ is a known outcome obtained from the product of outcomes of measurements in $V$. As $u\in\langle C\rangle$, $u$ has an associated stabilizer $s_{u}\in\langle S_{0}\rangle$. If $u=s_{u}$, then $u\cdot s_{u}=I$ has known outcome $+1$. To show that the outcome of $u\cdot s_{u}$ is known, we first show that for any element $c\in C$ with its associated stabilizer $s_{c}$, $c\cdot s_{c}$ always has a known outcome denoted by $O(c\cdot s_{c})$. We prove this by induction. Suppose the statement is true for all elements in $C$ before the $k^{\mathrm{th}}$ measurements. Then if we measure $m$ as the $(k+1)^{\mathrm{th}}$ measurement, an element $c\in C$ can be updated using one of the four cases in Step 2 of the distance algorithm. In all 4 cases, the statement remains true. Case 1 and Case 3 do not update $C$, so the statement is trivially true in these 2 cases. In Case 2, if $c_{1}$ has an associated stabilizer $s_{c_{1}}$ and $O(c_{1}\cdot s_{c_{1}})$ is known and $c_{2}$ has an associated stabilizer $s_{c_{2}}$ and $O(c_{2}\cdot s_{c_{2}})$ is known by the induction hypothesis. Then, if applying the update rule removes $c_{2}$ and updates $c_{1}\to c_{1}\cdot c_{2}$, the associated stabilizer is updated to $s_{1}\cdot s_{2}$, and the outcome $O((c_{1}\cdot c_{2})\cdot(s_{c_{1}}\cdot s_{c_{2}}))$ is given by $O(c_{1}\cdot s_{c_{1}})\cdot O(c_{2}\cdot s_{c_{2}})$ which is known. In Case 4, if an element $c$ in $C$, with an associated stabilizer $s_{c}$, is updated by the measurement $m\in V$ such that $c\to c\cdot m$, then the associated stabilizer remains the same as $s_{c}$. The outcome $O((c\cdot m)\cdot s_{c})$ is given by $O(c\cdot s_{c})\cdot O(m)$ which is known. Thus, we showed that for any element $c\in C$ with an associated stabilizer $s_{c}\in\langle S_{0}\rangle$, $O(c\cdot s_{c})$ is always known. Since $O(s_{u})=O(u\cdot s_{u})\cdot O(u)$, and $O(u\cdot s_{u})$ is a known outcome by the argument above, any element in $\langle C\rangle\cap\langle V\rangle$ gives an unmasked stabilizer generator $s_{u}\in\langle S_{0}\rangle$ along with its syndrome $O(s_{u})$. ∎
@@
We prove this by considering two choices of generators for $C$ and $V$, with the elements in the two choices for $C$ differing by some elements in $\langle V\rangle$. This is necessary since Case 4 of the distance algorithm will remove an arbitrary anti-commuting element from $V$ and multiply it to some elements in $C$. We show that for both choices, we obtain the same set of unmasked stabilizers. One property that we will make use of heavily here is that the stabilizer update rules in Lemma 3 are basis independent. Let $V_{1},V_{2}$ be two arbitrary bases for $\langle V\rangle$ such that $\langle V_{1}\rangle=\langle V_{2}\rangle=\langle V\rangle$. Let $C_{1},C_{2}$ be two sets for $C$. Elements in $C_{1}$ can be expressed in terms of elements in $\langle C_{2}\rangle$ and some generators $\{v_{1},\cdots,v_{k}\}\subset\langle V\rangle$ and vice versa. Note that for $C_{1}$ and $C_{2}$, we must be careful that there may be elements that have the same Pauli operator but different associated stabilizers, thus we require that in our choices the groups formed by the associated stabilizers of the elements in the two sets must be equal, that is for any element $c_{1}\in\langle C_{1}\rangle$, $\exists c_{2}\in\langle C_{2}\rangle$ with the same associated stabilizer generator as $c_{1}$ such that $c_{1}=c_{2}\cdot v$, with $v\in\langle V\rangle$, and vice versa. Initially, both choices of generators are equivalent in finding the unmasked stabilizers. Next, we show that this holds true after an arbitrary measurement $m$, by checking for all 4 cases of the commutation relations on $C_{1},V_{1}$. We first check for Case 1 and Case 2. Then, we check the update when it is Case 3 for both $C_{1},V_{1}$ and $C_{2},V_{2}$. Lastly, we check the update when it is Case 3 for $C_{1},V_{1}$ and either Case 3 or 4 for $C_{2},V_{2}$ without loss of generality.

### v2

@@ 4.1 Definition of Distance for a Dynamical Code
Figure 1: The figure illustrates an example of a Floquet code which is a special case of dynamical codes consisting of periodic measurements (In this case, 4 rounds of measurements). Each round of measurements consists of commuting Paulis. The $i^{\mathrm{th}}$ round of measurements brings the code from the ISG defined by $S_{i-1}$ to the next ISG defined by $S_{i(\mathrm{mod}\>4)}$.
@@
We begin by outlining the problem under consideration. For a given dynamical code, what is the set of errors that an ISG can correct given the syndrome information acquired through subsequent rounds of measurements? Using this information, what is the distance of the ISG? Without loss of generality, we examine the ISG at round 0 of the code and assume that all syndromes for this ISG come from the measurement sequence.
When defining the distance, we assume that the errors only occur at round 0 of the code and we also assume that if the errors are trivial, we know what the syndromes for the stabilizers in round 0 will be. The assumption that all errors occur in round 0 means that this notion of distance does not fully capture fault tolerance of the dynamical code. Instead, it is the analogue to the distance of a static quantum error-correcting code, which gives a necessary condition for fault tolerance, but not a sufficient one. However, note that the distance algorithm has wider applicability: It works by classifying the generators of the ISG as unmasked, temporarily masked, and permanently masked, which makes no reference to the actual errors in the system. Only the definition of distance itself uses the assumption that the errors are at round 0; the classification of stabilizer generators produced by the algorithm is still correct when errors can occur at any point.
@@
Another example is the subsystem surface codes with three-qubit check operators in [32]. In this case, $d_{\mathrm{subsystem}}\leq d_{\mathrm{ISG}}$. If the gauges are fixed in the x basis, then the distance for a $3L^{2},2,L]]$ code is given by $L$ for $X$ logical and $2L$ for $Z$ logical, but for $d_{\mathrm{subsystem}}$, the minimum weight of dressed logical is $L$ for both $X$ and $Z$ logical operators.
@@ 4.2 Masking and Unmasked Distance of a Dynamical Code
One can extend the definition of $T$ to that of a group: Define $\tilde{T}=\langle T\rangle U$. Then, $\tilde{T}$ consists of the group of unmasked and temporarily masked stabilizers. We have $U\subseteq\tilde{T}$ and $\T~P=\langle S\rangle\backslash\tilde${T}.
@@
However, there are cases where even though two errors differ on the syndrome for masked generators, they do not affect our measurements of logical states. Let’s assume $E^{\dagger}F\in\mathcal{N}(U)\backslash\mathcal{N}(S)$, so $E$ and $F$ have the same unmasked syndrome but different masked syndromes. If $L$ and $L^{\prime}$ are two representatives for the same logical qubit, then measuring either is normally a valid way of measuring the logical value. Now, if $E^{\dagger}F,L]=0$, then it does not matter whether we have an error $E$ or an error $F$; in either case, we will get the same eigenvalue for $L$ when we measure it. On the other hand, if $\{E^{\dagger}F,L^{\prime}\}=0$, then the error $E$ followed by measuring $L^{\prime}$ will give us a different result than the error $F$ followed by measuring $L^{\prime}$. In other words, we can pick a coset of logical representatives for which $E^{\dagger}F,L]=0$ and this will allow us to correctly deduce the logical outcomes despite a larger set of errors that now consists of indistinguishable errors like $E$ and $F$.
@@
The distance is defined as the minimum weight of $E^{\dagger}F$ that does not satisfy Eq. (8), which is the complement of the set $(P_{\mathrm{n}}-\mathcal{N}(U))\cup\langle S\rangle]\cdot\mathcal{G}$. This is exactly given by
@@
With the fixed choices of destabilizers, we have completely determined the set of bare logical operators, up to $U$. We see that there are 3 equivalent notions here: Choosing a gauge group $\mathcal{G}$ is the same as fixing the set of destabilizers which is also equivalent to choosing the bare logical representatives. We can see that doing so allows us to correct for more errors, since it also allows for errors that are destabilizers of the masked stabilizers, because these errors satisfy the condition that $E^{\dagger}F,L]=0$, for all bare logical representatives $L$s of the new gauge group.
@@ 4.3 Logical Operators of a Dynamical Code
Suppose we initialize a dynamical code that consists of $l$ rounds of measurements in an eigenstate of $L_{0}$, which is fixed as our choice of logical representative for the $0^{\mathrm{th}}$ ISG, with the notations illustrated in Figure 3. Each round of measurements consists of commuting Pauli operators which form the set $M_{i}$ for the $i^{\mathrm{th}}$ round. Let $E_{k}$ be the net error from the time after the $k^{\mathrm{th}}$ round of measurements till the final round of the code. Then, $E_{0}$ is the net error since initialization till the final round of the code. The Pauli operators $L_{0}$ and $L_{l}$ are related by $L_{l}=L_{0}\cdot s_{0}\cdot\prod_{i=1}^{l-1}m_{i}$ up to $\pm 1$ sign difference. $m_{i}$ is an element of $\langle M_{i}\rangle$ and $O(m_{i})$ is the outcome of $m_{i}$ which can be obtained through the outcomes of elements in $M_{i}$. $s_{0}$ is an element of $\langle S_{0}\rangle$ which is the stabilizer group for round 0 of the dynamical code and $O(s_{0})$ is the outcome of $s_{0}$. $A\odot B$ is the binary symplectic product of A with B as defined in Subsection 3.3.
@@
The case where errors only occurred in round 0 follows as a special case.
@@
[…]
ISG Stabilizers Logical Events
@@ 22 $s_{1},s_{2},\cdots$ $L_{1}$
33 $s_{1},s_{3},\cdots$ $L_{1}\cdot s_{2}$
Assuming $L_{1},e]=0$, even though $L_{1}\cdot s_{2}$ anti-commutes with $e$, the correct outcome of the logical operator is just given by $O(L_{1})\cdot O(s_{2})$, independent of the error $e$ that happened earlier in the measurement schedule.
[…]
@@ 33 $s_{1},s_{3},\cdots$ $L_{1}\cdot s_{2}$
On the other hand, if the error $e$ occurred in round 2, after measuring $s_{2}$, we see a change in the logical outcome. Here, we again assume that $L_{1},e]=0$
ISG Stabilizers Logical Events
11 $s_{1},\cdots$ $L_{1}$
@@
33 $s_{1},s_{3},\cdots$ $L_{1}\cdot s_{2}$
[…]
@@ 4.4 Syndromes through Measurements in a Dynamical Code
$e_{0}\odot\prod_{i=1}^{f}m_{i}+e_{1}\odot\prod_{i=2}^{f}m_{i}+\cdots+e_{f-1}\odot m_{f}\mod\>2$ (14)
$=$ $\prod_{j}\left(e_{j}\odot\prod_{i=j+1}^{f}m_{i}\right)\mod\>2$ (15)
@@
As the decoding problem is much harder for the situation where errors occur in all rounds of the dynamical code and can depend on the details of the dynamical code, we consider only the simpler scenario with errors happening only in round 0 of the dynamical code, so that the outcomes for stabilizers in round 0 correspond only to syndromes of errors in round 0. From here, we can obtain an upper bound in the distance of the code.
@@ 5 The Distance Algorithm
We consider a dynamical code $\mathcal{Q}$ with $l$ rounds of measurements. The ISGs are denoted by $S_{0},S_{1},\cdots,S_{l}$ starting with the ISG $S_{0}$ in round 0. The sets of commuting measurements in each round are denoted by $M_{1},M_{2},\cdots$, where $M_{i}$ is the set of measurements made after the $(i-1)^{\mathrm{th}}$ ISG.
The $l$ rounds of measurements can also be interpreted as the window of measurements for obtaining the syndrome information to decode for errors in round 0. We assume that some errors occurred in round 0 and no more errors happen in the later rounds of measurements and we only use $l$ rounds of measurements as our measurement window for learning the syndrome bits in round 0. To gain a better understanding of the distance of the ISG in round 0, we need to know for each syndrome bit in round 0, which of the following 3 cases does it fall under given only the first $l$ rounds of measurements:
@@ 5.3 The Distance Algorithm
      Case 1: $c,m]=[v,m]=0,\forall c\in C,\forall v\in V$ Update rule: Add $m$ to $V$ if $m\notin\langle V\rangle$ and record down $O(m)$.
@@
      Case 3: $c,m]=\{v,m\}=0,\forall c\in C$ and for some $v\in V$
@@
[…]
@@ 5.3.1 Unmasked Stabilizers and their Syndromes
Here, $\alpha_{i},\beta_{i}]$ is a non-zero 2n-dimensional vector and $\bullet$ stands for an arbitrary n-dimensional vector.
The generators of $\langle C\rangle\cap\langle V\rangle$ are given by the set $\{F^{-1}([e_{i},f_{i}])\},1\leq i\leq j$ and the identities from the 0 vector rows of Eq. (17), as each 0 vector corresponds to a different associated stabilizer generator.
@@ 5.3.2 Permanently Masked Stabilizers and their Destabilizers
    Initialize $L$ as an empty set to keep track of the gauge operators for $\mathcal{Q}_{2}$ .
@@ 5.4 Examples for the Distance Algorithm
Round $C$ $V$
0 $x_{1}x_{2}x_{3}x_{4}x_{5}x_{6}$
11 $x_{1}x_{2}x_{3}x_{4}x_{5}x_{6}$ $x_{5}x_{6}$
22 $x_{1}x_{2}x_{3}x_{4}$ $z_{6}z_{7}$
33 $x_{1}x_{2}x_{3}x_{4}$ $z_{6}z_{7},x_{1}x_{2},x_{3}x_{4}$
$\langle C\rangle\cap\langle V\rangle=\{x_{1}x_{2}x_{3}x_{4}\}$
Using algorithm 2, we can substitute into the equation $O(s_{j})=O(u_{j})\cdot O(u_{j}\cdot s_{j})$ with $O(u_{j})=O(x_{1}x_{2}x_{3}x_{4})$ and $O(u_{j}\cdot s_{j})=O(x_{5}x_{6})$. This gives the outcome for $x_{1}x_{2}x_{3}x_{4}x_{5}x_{6}$ which is an unmasked stabilizer for this measurement sequence.
@@ 5.5 Proof for the Distance Algorithm
Let $u$ be an element in $\tilde{U}$ and let its associated stabilizer be $s_{u}$. We prove that both $O(u)$ and $O(u\cdot s_{u})$ are known. This shows that $O(s_{u})$ has a known syndrome given by $O(u)\cdot O(u\cdot s_{u})$.
Since $u\in\langle V\rangle$, $O(u)$ is a known outcome obtained from the product of outcomes of measurements in $V$. As $\in\langle C\rangle, uu$ has an associated stabilizer $s_{u}\in\langle S_{0}\rangle$. If $u=s_{u}$, then $u\cdot s_{u}=I$ has known outcome +1+1.
To show that the outcome of $u\cdot s_{u}$ is known, we first show that for any element $c\in C$ with its associated stabilizer $s_{c}$, $c\cdot s_{c}$ always has a known outcome denoted by $O(c\cdot s_{c})$. We prove this by induction.
Suppose the statement is true for all elements in $C$ before the $k^{\mathrm{th}}$ measurements.
Then if we measure $m$ as the $(k+1)^{\mathrm{th}}$ measurement, an element $c\in C$ can be updated using one of the four cases in Step 2 of the distance algorithm. In all 4 cases, the statement remains true.
Case 1 and Case 3 do not update $C$, so the statement is trivially true in these 2 cases.
In Case 2, if $c_{1}$ has an associated stabilizer $s_{c_{1}}$ and $O(c_{1}\cdot s_{c_{1}})$ is known and $c_{2}$ has an associated stabilizer $s_{c_{2}}$ and $O(c_{2}\cdot s_{c_{2}})$ is known by the induction hypothesis. Then, if applying the update rule removes $c_{2}$ and updates $c_{1}\to c_{1}\cdot c_{2}$, the associated stabilizer is updated to $s_{1}\cdot s_{2}$, and the outcome $O((c_{1}\cdot c_{2})\cdot(s_{c_{1}}\cdot s_{c_{2}}))$ is given by $O(c_{1}\cdot s_{c_{1}})\cdot O(c_{2}\cdot s_{c_{2}})$ which is known.
In Case 4, if an element $c$ in $C$, with an associated stabilizer $s_{c}$, is updated by the measurement $m\in V$ such that $c\to c\cdot m$, then the associated stabilizer remains the same as $s_{c}$. The outcome $O((c\cdot m)\cdot s_{c})$ is given by $O(c\cdot s_{c})\cdot O(m)$ which is known.
Thus, we showed that for any element $c\in C$ with an associated stabilizer $s_{c}\in\langle S_{0}\rangle$, $O(c\cdot s_{c})$ is always known.
Since $O(s_{u})=O(u\cdot s_{u})\cdot O(u)$, and $O(u\cdot s_{u})$ is a known outcome by the argument above, any element in $\langle C\rangle\cap\langle V\rangle$ gives an unmasked stabilizer generator $s_{u}\in\langle S_{0}\rangle$ along with its syndrome $O(s_{u})$.
@@
We prove this by considering two choices of generators for $C$ and $V$, with the elements in the two choices for $C$ differing by some elements in $\langle V\rangle$. This is necessary since Case 4 of the distance algorithm will remove an arbitrary anti-commuting element from $V$ and multiply it to some elements in $C$. We show that for both choices, we obtain the same set of unmasked stabilizers. One property that we will make use of heavily here is that the stabilizer update rules in Lemma 3 are basis independent.
Let $V_{1},V_{2}$ be two arbitrary bases for $\langle V\rangle$ such that $\langle V_{1}\rangle=\langle V_{2}\rangle=\langle V\rangle$. Let $C_{1},C_{2}$ be two sets for $C$. Elements in $C_{1}$ can be expressed in terms of elements in $\langle C_{2}\rangle$ and some generators $\{v_{1},\cdots,v_{k}\}\subset\langle V\rangle$ and vice versa. Note that for $C_{1}$ and $C_{2}$, we must be careful that there may be elements that have the same Pauli operator but different associated stabilizers, thus we require that in our choices the groups formed by the associated stabilizers of the elements in the two sets must be equal, that is for any element $c_{1}\in\langle C_{1}\rangle$, $\exists c_{2}\in\langle C_{2}\rangle$ with the same associated stabilizer generator as $c_{1}$ such that $c_{1}=c_{2}\cdot v$, with $v\in\langle V\rangle$, and vice versa.
Initially, both choices of generators are equivalent in finding the unmasked stabilizers. Next, we show that this holds true after an arbitrary measurement $m$, by checking for all 4 cases of the commutation relations on $C_{1},V_{1}$. We first check for Case 1 and Case 2. Then, we check the update when it is Case 3 for both $C_{1},V_{1}$ and $C_{2},V_{2}$. Lastly, we check the update when it is Case 3 for $C_{1},V_{1}$ and either Case 3 or 4 for $C_{2},V_{2}$ without loss of generality.